# Is Artificial Intelligence Generated Image Detection a Solved Problem?

### v1

@@ 1 Introduction
Recent advancements in generative models, such as GANs [1, 2, 3, 4, 5, 6] and diffusion models [7, 8, 9, 10], have demonstrated remarkable capabilities in synthesizing high-quality images that closely resemble real-world scenes. While these technologies have enabled various applications, including personalized portraits [11], virtual try-on, and content creation [12, 13], they also raise significant concerns regarding the authenticity of visual content and its potential misuse in misinformation dissemination [14], deepfake generation [15], and copyright infringement [16, 8].
To address these concerns, Artificial Intelligence Generated Image (AIGI) detection technologies have been developed to differentiate synthetic images from real ones. Most existing approaches [17, 18, 19, 20, 21, 22] rely on training binary classifiers to distinguish between real and generated content, with some studies reporting near-perfect performance. In light of reported detection accuracies exceeding 95%, a critical question emerges: Is Artificial Intelligence Generated Image detection a solved problem?
To answer this question, we propose a novel and comprehensive benchmark, termed AIGIBench, for AIGI detection. As illustrated in Figure 2 of the Sec. A.1, we present an end-to-end pipeline designed to assess AIGI detection in real-world scenarios. During the training phase, a diverse collection of real and synthetic images is assembled, and various data augmentation techniques are applied to improve detector robustness. In the testing phase, the detector is evaluated on its ability to identify the authenticity of images originating from unknown sources and subjected to unknown degradations. Prior to detection, each image undergoes pre-processing, such as cropping or resizing to ensure compatibility with the trained detector. The detector then outputs a binary classification indicating whether the image is real or AI-generated. Accordingly, to systematically evaluate the real-world performance of state-of-the-art AIGI detectors, AIGIBench defines four core tasks that mirror practical challenges often overlooked in idealized test environments: i) Generalization Assessment: Multi-source, ii) Robustness Assessment: Multi-degradation, iii) Data Augmentation Variation Assessment: Identifying the Most Effective Augmentation, and iv) Test Data Pre-processing Assessment: Identifying the Most Effective Pre-processing. These task variations reflect common challenges in practical applications but are often overlooked in idealized test environments where most existing detectors are developed and evaluated.
@@
Benchmark $\downarrow$ Dataset $\rightarrow$ Generative Methods ~2022 2023 2024~ General Content GAN & Diffusion Image-based & Noise-based Social Networks AI-painting Communities
@@
Benchmark $\downarrow$ Evaluation $\rightarrow$ Detection Methods ~2022 2023 2024~ Generalization Robust Data Augmentation Test Data Processing
@@ 2.2 Evaluation Datasets
Fake Image Collection. To ensure the quality and fairness of our dataset, we collect substantially more images than necessary through both generator and web API, and applied the following processing pipeline: i) To mitigate the presence of visually similar images, we use CLIP [55] to extract image features and remove highly similar instances, promoting diversity and fairness in detection; ii) To enhance image quality and better reflect real-world social environments, we employ the CLIP-based aesthetic score predictor [56] to evaluate image aesthetics and discarded images with low scores; iii) Finally, manual screening is conducted to remove obviously fake images, further increasing the dataset’s realism and challenge. We collect diverse fake images from various sources and organize them into five levels based on the generation pipelines of different models.
@@
ii) Diffusion for Text-to-Image Generation: Given the rapid advancements in diffusion models, our benchmark includes seven representative models: SD-XL [36], SD-3 [37], DALLE-3 [38], Midjourney-v6 [39], FLUX.1-dev [40], Imagen-3 [41], and GLIDE [42]. These text-to-image diffusion models generate images by iteratively denoising random noise under the guidance of textual prompts, making the quality and diversity of input text crucial for dataset construction. To this end, we leverage the Gemini [57] API with carefully designed prompts to synthesize diverse and high-quality image descriptions. For instance, a prompt for generating realistic human descriptions might read: “To create varied descriptions of people in photo or realistic style, I need 1000 distinct sentences, each 20–25 words, please help me. Example: ‘woman wearing a red dress in the park, Disney cartoon style.’”. We generate 1500 sentences for each of four categories—people, animals, objects, and landscapes—to ensure broad coverage and content diversity.
@@ 3 Experiments and Discussion
3.1 Task 1: Generalization Assessment
3.2 Task 2: Robustness Assessment
@@ 3.3 Task 3: Data Augmentation Variation Assessment
Detectors $\rightarrow$ CLIPDetection FreqNet NPR DFFreq LaDeDa SAFE
@@ Appendix A Appendix
A.1 Pipeline of AIGI Detection in Real-World Scenarios
Figure 2 illustrates the end-to-end pipeline for AIGI detection in real-world scenarios. In the training phase, a diverse collection of real and generated images is assembled, and various data augmentation strategies are applied to improve the detector’s robustness. During the testing phase, the detector evaluates the authenticity of images originating from unknown sources and potentially affected by unknown degradations. Before inference, each image undergoes pre-processing, such as cropping or resizing, to ensure compatibility with the detector. The final output is a classification result indicating whether the image is authentic or generated.
@@ A.1 Pipeline of AIGI Detection in Real-World Scenarios
A.2 More Results on Task 1: Generalization Assessment
@@
A.3 More Results on Task 2: Robustness Assessment
@@
A.4 More Results on Task 3: Data Augmentation Variation Assessment
@@
A.5 More Results on Task 4: Test Data Pre-processing Assessment

### v2

@@ 1 Introduction
Recent advancements in generative models, such as GANs [1, 2, 3, 4, 5, 6] and diffusion models [7, 8, 9, 10], have demonstrated remarkable capabilities in synthesizing high-quality images that closely resemble real-world scenes. While these technologies have enabled various applications, including personalized portraits [11], virtual try-on, and content creation [12, 13], they also raise significant concerns regarding the authenticity of visual content and its potential misuse in misinformation dissemination [14, 15], deepfake generation [16], and copyright infringement [17, 8].
To address these concerns, Artificial Intelligence Generated Image (AIGI) detection technologies have been developed to differentiate synthetic images from real ones. Most existing approaches [18, 19, 20, 21, 22, 23, 24, 25, 26, 27, 28] rely on training binary classifiers to distinguish between real and generated content, with some studies reporting near-perfect performance. In light of reported detection accuracies exceeding 95%, a critical question emerges: Is Artificial Intelligence Generated Image detection a solved problem?
To answer this question, we propose a novel and comprehensive benchmark, termed AIGIBench, for AIGI detection. As illustrated in Figure 1, we present an end-to-end pipeline designed to assess AIGI detection in real-world scenarios. During the training phase, a diverse collection of real and synthetic images is assembled, and various data augmentation techniques are applied to improve detector robustness. In the testing phase, the detector is evaluated on its ability to identify the authenticity of images originating from unknown sources and subjected to unknown degradations. Prior to detection, each image undergoes pre-processing, such as cropping or resizing to ensure compatibility with the trained detector. The detector then outputs a binary classification indicating whether the image is real or AI-generated. Accordingly, to systematically evaluate the real-world performance of state-of-the-art AIGI detectors, AIGIBench defines four core tasks that mirror practical challenges often overlooked in idealized test environments: i) Generalization Assessment: Multi-source, ii) Robustness Assessment: Multi-degradation, iii) Data Augmentation Variation Assessment: Identifying the Most Effective Augmentation, and iv) Test Data Pre-processing Assessment: Identifying the Most Effective Pre-processing. These task variations reflect common challenges in practical applications but are often overlooked in idealized test environments where most existing detectors are developed and evaluated.
@@
Benchmark $\downarrow$ Dataset $\rightarrow$
Generative
Methods
~2022 2023 2024~
General
Content
GAN &
Diffusion
Image-based &
Noise-based
Social
Networks
AI-painting
Communities
@@ Methods
Benchmark $\downarrow$ Evaluation $\rightarrow$
Detection
Methods
~2022 2023 2024~ Generalization Robust
Data
Augmentation
Test Data
Processing
@@ 2.2 Evaluation Datasets
Fake Image Collection. To ensure the quality and fairness of our dataset, we collect substantially more images than necessary through both generator and web API, and applied the following processing pipeline: i) To mitigate the presence of visually similar images, we use CLIP [61] image embeddings with a cosine similarity threshold of 0.98 to identify and remove near-duplicate instances, promoting diversity and fairness in detection. More details are shown in Sec. A.5 of the Apendix; ii) To enhance image quality and better reflect real-world social environments, we employ the CLIP-based aesthetic score predictor [62] to evaluate image aesthetics and discarded images with low scores; iii) Finally, manual screening is conducted to remove obviously fake images, further increasing the dataset’s realism and challenge. We collect diverse fake images from various sources and organize them into five levels based on the generation pipelines of different models.
@@
ii) Diffusion for Text-to-Image Generation: Given the rapid advancements in diffusion models, our benchmark includes seven representative models: SD-XL [42], SD-3 [43], DALLE-3 [44], Midjourney-v6 [45], FLUX.1-dev [46], Imagen-3 [47], and GLIDE [48]. These text-to-image diffusion models generate images by iteratively denoising random noise under the guidance of textual prompts, making the quality and diversity of input text crucial for dataset construction. To this end, we leverage the Gemini [63] API with carefully designed prompts to synthesize diverse and high-quality image descriptions. For instance, a prompt for generating realistic human descriptions might read: "To create varied descriptions of people in photo or realistic style, I need 1000 distinct sentences, each 20–25 words, please help me. Example: ’woman wearing a red dress in the park, Disney cartoon style.’". We generate 1500 sentences for each of four categories—people, animals, objects, and landscapes—to ensure broad coverage and content diversity.
@@ 3.1 Task 1: Generalization Assessment
Detectors $\rightarrow$ CLIPDetection FreqNet NPR DFFreq LaDeDa SAFE
3.2 Task 2: Robustness Assessment
@@ 4 Conclusions, Limitations, and Board Impacts
5 Acknowledgments
This work was supported the National Natural Science Foundation of China under grant U22B2062, 62502215, 62172233, 62472231, 62172232, 62402229; the Natural Science Foundation of Jiangsu under Grants BK20250735; the Jiangsu Provincial Science and Technology Major Project (No. BG2024042); the General Program of Natural Science Research in Universities of Jiangsu under Grants 25KJB520027.
@@ Appendix A Appendix
A.1 More Results on Task 1: Generalization Assessment
@@
A.2 More Results on Task 2: Robustness Assessment
@@
A.3 More Results on Task 3: Data Augmentation Variation Assessment
@@
A.4 More Results on Task 4: Test Data Pre-processing Assessment
@@
Table 17: F.Acc results on the original unfiltered dataset (main score) and the curated dataset (subscript indicates the difference). A positive subscript (e.g., +2.9) indicates that the detector performed worse on the curated dataset, implying that the curated dataset is more challenging.
Detector FaceSwap IP-Adapter Midjourney SD3 StyleSwim
CNNDetection $1.8_{+0.4}$ $5.0_{-1.0}$ $6.1_{+0.3}$ $16.2_{+2.9}$ $8.7_{+1.8}$
CLIPDetection $29.8_{+2.5}$ $94.3_{+2.3}$ $78.5_{-2.1}$ $92.1_{+1.5}$ $97.9_{-0.2}$
DFFreq $0.9_{+0.6}$ $76.8_{-1.3}$ $55.2_{+1.2}$ $75.7_{+2.3}$ $87.9_{+7.1}$
Ladeda $0.7_{+0.7}$ $92.6_{+2.0}$ $84.6_{+1.2}$ $98.5_{-0.5}$ $96.3_{-1.0}$
AIDE $12.0_{-2.3}$ $94.1_{+0.6}$ $80.1_{+0.3}$ $99.3_{+0.0}$ $91.4_{+9.4}$
SAFE $22.9_{+19.6}$ $99.3_{+9.5}$ $98.5_{+1.3}$ $86.3_{-5.4}$ $99.9_{+0.6}$
@@
A.5 More Confirmatory Analyses
In our data curation process, we employed CLIP image embeddings with a cosine similarity threshold of 0.98 to identify and remove near-duplicate instances. Additionally, we used an off-the-shelf aesthetic quality predictor and retained only high-resolution (4K) images with the highest aesthetic scores to increase the difficulty and discriminative power of the detection benchmark. Low-quality or visually homogeneous images often reduce the evaluation challenge and may artificially inflate detector performance. By removing redundant and low-aesthetic samples, we ensure that the curated dataset better reflects real-world, high-fidelity scenarios and enables a more rigorous and meaningful evaluation of detection capabilities.
To further verify its effectiveness, Table 17 presents the F.Acc (Fake Accuracy) results of six representative detectors on both the curated dataset and the original unfiltered dataset. The main value in each cell corresponds to the detector’s performance on the original unfiltered dataset, which indicates the absolute performance difference between the unfiltered and curated datasets. A positive subscript (e.g. +2.9) indicates that the detector performed worse on the curated dataset, implying that the curated dataset is more challenging. Conversely, a negative subscript (e.g. -1.0) suggests improved performance. Across most models and subsets, we observe a consistent decline in F.Acc following curation, particularly for state-of-the-art detectors such as AIDE and SAFE. These results underscore the increased challenge presented by the curated dataset and validate our filtering approach as a means of constructing a more robust benchmark.
@@ A.7 Details of Detection Models Evaluated in Our AIGIBench
NeurIPS Paper Checklist
  1. 1.
    Claims
  2. Question: Do the main claims made in the abstract and introduction accurately reflect the paper’s contributions and scope?
  3. Answer: [Yes]
  4. Justification: Our contributions are briefly summarized within the "introduction" section
  5. Guidelines:
    • The answer NA means that the abstract and introduction do not include the claims made in the paper.
    • The abstract and/or introduction should clearly state the claims made, including the contributions made in the paper and important assumptions and limitations. A No or NA answer to this question will not be perceived well by the reviewers.
    • The claims made should match theoretical and experimental results, and reflect how much the results can be expected to generalize to other settings.
    • It is fine to include aspirational goals as motivation as long as it is clear that these goals are not attained by the paper.
  6. 2.
    Limitations
  7. Question: Does the paper discuss the limitations of the work performed by the authors?
  8. Answer: [Yes]
  9. Justification: We create a subsection "Limitation and Future Works" in our paper. AIGIBench currently considers two widely adopted training settings: i) training on ProGAN, and ii) training on ProGAN combined with SD-v1.4.
  10. Guidelines:
    • The answer NA means that the paper has no limitation while the answer No means that the paper has limitations, but those are not discussed in the paper.
    • The authors are encouraged to create a separate "Limitations" section in their paper.
    • The paper should point out any strong assumptions and how robust the results are to violations of these assumptions (e.g., independence assumptions, noiseless settings, model well-specification, asymptotic approximations only holding locally). The authors should reflect on how these assumptions might be violated in practice and what the implications would be.
    • The authors should reflect on the scope of the claims made, e.g., if the approach was only tested on a few datasets or with a few runs. In general, empirical results often depend on implicit assumptions, which should be articulated.
    • The authors should reflect on the factors that influence the performance of the approach. For example, a facial recognition algorithm may perform poorly when image resolution is low or images are taken in low lighting. Or a speech-to-text system might not be used reliably to provide closed captions for online lectures because it fails to handle technical jargon.
    • The authors should discuss the computational efficiency of the proposed algorithms and how they scale with dataset size.
    • If applicable, the authors should discuss possible limitations of their approach to address problems of privacy and fairness.
    • While the authors might fear that complete honesty about limitations might be used by reviewers as grounds for rejection, a worse outcome might be that reviewers discover limitations that aren’t acknowledged in the paper. The authors should use their best judgment and recognize that individual actions in favor of transparency play an important role in developing norms that preserve the integrity of the community. Reviewers will be specifically instructed to not penalize honesty concerning limitations.
  11. 3.
    Theory assumptions and proofs
  12. Question: For each theoretical result, does the paper provide the full set of assumptions and a complete (and correct) proof?
  13. Answer: [N/A]
  14. Justification: [N/A]
  15. Guidelines:
    • The answer NA means that the paper does not include theoretical results.
    • All the theorems, formulas, and proofs in the paper should be numbered and cross-referenced.
    • All assumptions should be clearly stated or referenced in the statement of any theorems.
    • The proofs can either appear in the main paper or the supplemental material, but if they appear in the supplemental material, the authors are encouraged to provide a short proof sketch to provide intuition.
    • Inversely, any informal proof provided in the core of the paper should be complemented by formal proofs provided in appendix or supplemental material.
    • Theorems and Lemmas that the proof relies upon should be properly referenced.
  16. 4.
    Experimental result reproducibility
  17. Question: Does the paper fully disclose all the information needed to reproduce the main experimental results of the paper to the extent that it affects the main claims and/or conclusions of the paper (regardless of whether the code and data are provided or not)?
  18. Answer: [Yes]
  19. Justification: See the Appendix for details on the adopted dataset and detection methods.
  20. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • If the paper includes experiments, a No answer to this question will not be perceived well by the reviewers: Making the paper reproducible is important, regardless of whether the code and data are provided or not.
    • If the contribution is a dataset and/or model, the authors should describe the steps taken to make their results reproducible or verifiable.
    • Depending on the contribution, reproducibility can be accomplished in various ways. For example, if the contribution is a novel architecture, describing the architecture fully might suffice, or if the contribution is a specific model and empirical evaluation, it may be necessary to either make it possible for others to replicate the model with the same dataset, or provide access to the model. In general. releasing code and data is often one good way to accomplish this, but reproducibility can also be provided via detailed instructions for how to replicate the results, access to a hosted model (e.g., in the case of a large language model), releasing of a model checkpoint, or other means that are appropriate to the research performed.
    • While NeurIPS does not require releasing code, the conference does require all submissions to provide some reasonable avenue for reproducibility, which may depend on the nature of the contribution. For example
      1. (a)
        If the contribution is primarily a new algorithm, the paper should make it clear how to reproduce that algorithm.
      2. (b)
        If the contribution is primarily a new model architecture, the paper should describe the architecture clearly and fully.
      3. (c)
        If the contribution is a new model (e.g., a large language model), then there should either be a way to access this model for reproducing the results or a way to reproduce the model (e.g., with an open-source dataset or instructions for how to construct the dataset).
      4. (d)
        We recognize that reproducibility may be tricky in some cases, in which case authors are welcome to describe the particular way they provide for reproducibility. In the case of closed-source models, it may be that access to the model is limited in some way (e.g., to registered users), but it should be possible for other researchers to have some path to reproducing or verifying the results.
  21. 5.
    Open access to data and code
  22. Question: Does the paper provide open access to the data and code, with sufficient instructions to faithfully reproduce the main experimental results, as described in supplemental material?
  23. Answer: [Yes]
  24. Justification: All data and code are publicly available for full open access. Please ref to https://github.com/HorizonTEL/AIGIBench.
  25. Guidelines:
    • The answer NA means that paper does not include experiments requiring code.
    • Please see the NeurIPS code and data submission guidelines (https://nips.cc/public/guides/CodeSubmissionPolicy) for more details.
    • While we encourage the release of code and data, we understand that this might not be possible, so “No” is an acceptable answer. Papers cannot be rejected simply for not including code, unless this is central to the contribution (e.g., for a new open-source benchmark).
    • The instructions should contain the exact command and environment needed to run to reproduce the results. See the NeurIPS code and data submission guidelines (https://nips.cc/public/guides/CodeSubmissionPolicy) for more details.
    • The authors should provide instructions on data access and preparation, including how to access the raw data, preprocessed data, intermediate data, and generated data, etc.
    • The authors should provide scripts to reproduce all experimental results for the new proposed method and baselines. If only a subset of experiments are reproducible, they should state which ones are omitted from the script and why.
    • At submission time, to preserve anonymity, the authors should release anonymized versions (if applicable).
    • Providing as much information as possible in supplemental material (appended to the paper) is recommended, but including URLs to data and code is permitted.
  26. 6.
    Experimental setting/details
  27. Question: Does the paper specify all the training and test details (e.g., data splits, hyperparameters, how they were chosen, type of optimizer, etc.) necessary to understand the results?
  28. Answer: [Yes]
  29. Justification: The main experimental settings are presented in Section 2, with full implementation details provided in the Appendix.
  30. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • The experimental setting should be presented in the core of the paper to a level of detail that is necessary to appreciate the results and make sense of them.
    • The full details can be provided either with the code, in appendix, or as supplemental material.
  31. 7.
    Experiment statistical significance
  32. Question: Does the paper report error bars suitably and correctly defined or other appropriate information about the statistical significance of the experiments?
  33. Answer: [No]
  34. Justification: [No]
  35. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • The authors should answer "Yes" if the results are accompanied by error bars, confidence intervals, or statistical significance tests, at least for the experiments that support the main claims of the paper.
    • The factors of variability that the error bars are capturing should be clearly stated (for example, train/test split, initialization, random drawing of some parameter, or overall run with given experimental conditions).
    • The method for calculating the error bars should be explained (closed form formula, call to a library function, bootstrap, etc.)
    • The assumptions made should be given (e.g., Normally distributed errors).
    • It should be clear whether the error bar is the standard deviation or the standard error of the mean.
    • It is OK to report 1-sigma error bars, but one should state it. The authors should preferably report a 2-sigma error bar than state that they have a 96% CI, if the hypothesis of Normality of errors is not verified.
    • For asymmetric distributions, the authors should be careful not to show in tables or figures symmetric error bars that would yield results that are out of range (e.g. negative error rates).
    • If error bars are reported in tables or plots, The authors should explain in the text how they were calculated and reference the corresponding figures or tables in the text.
  36. 8.
    Experiments compute resources
  37. Question: For each experiment, does the paper provide sufficient information on the computer resources (type of compute workers, memory, time of execution) needed to reproduce the experiments?
  38. Answer: [No]
  39. Justification: [No]
  40. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • The paper should indicate the type of compute workers CPU or GPU, internal cluster, or cloud provider, including relevant memory and storage.
    • The paper should provide the amount of compute required for each of the individual experimental runs as well as estimate the total compute.
    • The paper should disclose whether the full research project required more compute than the experiments reported in the paper (e.g., preliminary or failed experiments that didn’t make it into the paper).
  41. 9.
    Code of ethics
  42. Question: Does the research conducted in the paper conform, in every respect, with the NeurIPS Code of Ethics https://neurips.cc/public/EthicsGuidelines?
  43. Answer: [Yes]
  44. Justification: [Yes]
  45. Guidelines:
    • The answer NA means that the authors have not reviewed the NeurIPS Code of Ethics.
    • If the authors answer No, they should explain the special circumstances that require a deviation from the Code of Ethics.
    • The authors should make sure to preserve anonymity (e.g., if there is a special consideration due to laws or regulations in their jurisdiction).
  46. 10.
    Broader impacts
  47. Question: Does the paper discuss both potential positive societal impacts and negative societal impacts of the work performed?
  48. Answer: [Yes]
  49. Justification: We create a subsection "Board Impacts." in our paper. AIGIBench enhances the detection of AI-generated images by offering a comprehensive and realistic evaluation framework. It exposes the limitations of current detectors and facilitates standardized comparisons, thereby driving the development of more robust and generalizable forensic tools. This work plays a critical role in addressing the societal risks posed by synthetic media, such as misinformation and digital fraud, and serves as a valuable resource for researchers, practitioners. However, a potential ethical concern arises from the possibility that malicious actors could misuse AIGIBench to improve the evasiveness of generative models.
  50. Guidelines:
    • The answer NA means that there is no societal impact of the work performed.
    • If the authors answer NA or No, they should explain why their work has no societal impact or why the paper does not address societal impact.
    • Examples of negative societal impacts include potential malicious or unintended uses (e.g., disinformation, generating fake profiles, surveillance), fairness considerations (e.g., deployment of technologies that could make decisions that unfairly impact specific groups), privacy considerations, and security considerations.
    • The conference expects that many papers will be foundational research and not tied to particular applications, let alone deployments. However, if there is a direct path to any negative applications, the authors should point it out. For example, it is legitimate to point out that an improvement in the quality of generative models could be used to generate deepfakes for disinformation. On the other hand, it is not needed to point out that a generic algorithm for optimizing neural networks could enable people to train models that generate Deepfakes faster.
    • The authors should consider possible harms that could arise when the technology is being used as intended and functioning correctly, harms that could arise when the technology is being used as intended but gives incorrect results, and harms following from (intentional or unintentional) misuse of the technology.
    • If there are negative societal impacts, the authors could also discuss possible mitigation strategies (e.g., gated release of models, providing defenses in addition to attacks, mechanisms for monitoring misuse, mechanisms to monitor how a system learns from feedback over time, improving the efficiency and accessibility of ML).
  51. 11.
    Safeguards
  52. Question: Does the paper describe safeguards that have been put in place for responsible release of data or models that have a high risk for misuse (e.g., pretrained language models, image generators, or scraped datasets)?
  53. Answer: [No]
  54. Justification: [No]
  55. Guidelines:
    • The answer NA means that the paper poses no such risks.
    • Released models that have a high risk for misuse or dual-use should be released with necessary safeguards to allow for controlled use of the model, for example by requiring that users adhere to usage guidelines or restrictions to access the model or implementing safety filters.
    • Datasets that have been scraped from the Internet could pose safety risks. The authors should describe how they avoided releasing unsafe images.
    • We recognize that providing effective safeguards is challenging, and many papers do not require this, but we encourage authors to take this into account and make a best faith effort.
  56. 12.
    Licenses for existing assets
  57. Question: Are the creators or original owners of assets (e.g., code, data, models), used in the paper, properly credited and are the license and terms of use explicitly mentioned and properly respected?
  58. Answer: [Yes]
  59. Justification: All images used in this paper were collected by the authors, with proper citations to the original sources where applicable.
  60. Guidelines:
    • The answer NA means that the paper does not use existing assets.
    • The authors should cite the original paper that produced the code package or dataset.
    • The authors should state which version of the asset is used and, if possible, include a URL.
    • The name of the license (e.g., CC-BY 4.0) should be included for each asset.
    • For scraped data from a particular source (e.g., website), the copyright and terms of service of that source should be provided.
    • If assets are released, the license, copyright information, and terms of use in the package should be provided. For popular datasets, paperswithcode.com/datasets has curated licenses for some datasets. Their licensing guide can help determine the license of a dataset.
    • For existing datasets that are re-packaged, both the original license and the license of the derived asset (if it has changed) should be provided.
    • If this information is not available online, the authors are encouraged to reach out to the asset’s creators.
  61. 13.
    New assets
  62. Question: Are new assets introduced in the paper well documented and is the documentation provided alongside the assets?
  63. Answer: [No]
  64. Justification: The new assets introduced in this paper consist of images generated using various generative models. We have publicly released these assets for community use.
  65. Guidelines:
    • The answer NA means that the paper does not release new assets.
    • Researchers should communicate the details of the dataset/code/model as part of their submissions via structured templates. This includes details about training, license, limitations, etc.
    • The paper should discuss whether and how consent was obtained from people whose asset is used.
    • At submission time, remember to anonymize your assets (if applicable). You can either create an anonymized URL or include an anonymized zip file.
  66. 14.
    Crowdsourcing and research with human subjects
  67. Question: For crowdsourcing experiments and research with human subjects, does the paper include the full text of instructions given to participants and screenshots, if applicable, as well as details about compensation (if any)?
  68. Answer: [N/A]
  69. Justification: [N/A]
  70. Guidelines:
    • The answer NA means that the paper does not involve crowdsourcing nor research with human subjects.
    • Including this information in the supplemental material is fine, but if the main contribution of the paper involves human subjects, then as much detail as possible should be included in the main paper.
    • According to the NeurIPS Code of Ethics, workers involved in data collection, curation, or other labor should be paid at least the minimum wage in the country of the data collector.
  71. 15.
    Institutional review board (IRB) approvals or equivalent for research with human subjects
  72. Question: Does the paper describe potential risks incurred by study participants, whether such risks were disclosed to the subjects, and whether Institutional Review Board (IRB) approvals (or an equivalent approval/review based on the requirements of your country or institution) were obtained?
  73. Answer: [N/A]
  74. Justification: [N/A]
  75. Guidelines:
    • The answer NA means that the paper does not involve crowdsourcing nor research with human subjects.
    • Depending on the country in which research is conducted, IRB approval (or equivalent) may be required for any human subjects research. If you obtained IRB approval, you should clearly state this in the paper.
    • We recognize that the procedures for this may vary significantly between institutions and locations, and we expect authors to adhere to the NeurIPS Code of Ethics and the guidelines for their institution.
    • For initial submissions, do not include any information that would break anonymity (if applicable), such as the institution conducting the review.
  76. 16.
    Declaration of LLM usage
  77. Question: Does the paper describe the usage of LLMs if it is an important, original, or non-standard component of the core methods in this research? Note that if the LLM is used only for writing, editing, or formatting purposes and does not impact the core methodology, scientific rigorousness, or originality of the research, declaration is not required.
  78. Answer: [Yes]
  79. Justification: We employ LLMs to generate descriptions for the synthesized images. The usage of LLMs is detailed in Section 2.
  80. Guidelines:
    • The answer NA means that the core method development in this research does not involve LLMs as any important, original, or non-standard components.
    • Please refer to our LLM policy (https://neurips.cc/Conferences/2025/LLM) for what should or should not be described.